### v4

@@ 5.2 Personalized Style Generation
TTF framework. In this field, StyleDrop [24] leverages adapter tuning to efficiently capture the style from a single reference image. This method demonstrates the effectiveness through iterative training, utilizing synthesized images refined by feedback mechanisms, like human evaluations and CLIP scores. This approach not only enhances style learning but also ensures that the generated styles align closely with human aesthetic judgments. Later, GAL [19] proposes an uncertainty-based evaluation strategy to filter high-quality synthetic-style data and uses a weighted schema to balance the contribution of the additional samples and the original reference. Furthermore, StyleAligned [99] focuses on maintaining stylistic consistency across a batch of images. This is achieved by using the first image as a reference, which acts as an additional key and value in the self-attention layers, ensuring that all subsequent images in the batch adhere to the same stylistic guidelines. Style-friendly [100] introduces a novel diffusion model fine-tuning approach that enhances personalized artistic style generation by adaptively biasing noise sampling toward higher noise levels, where stylistic features emerge.
PTA framework. For the PTA framework, StyleAdapter [82] employs a dual-path cross-attention mechanism within the PTA framework. This model introduces a specialized embedding module designed to extract and integrate global features from multiple style references. Diptych Prompting [101] utilizes an inpainting mechanism to draw another image with the same style of the reference part.
@@ 5.3 Personalized Face Generation
Personalized face generation aims to generate diverse identity images that adhere to text prompt specifications, utilizing only a few initial face images. Compared to general object personalization, the scope is narrowed to a specific class, humans. An obvious benefit is that one can readily leverage large-scale human-centric datasets [67, 102, 103] and utilize pre-trained models in well-developed areas, like face landmark detection [104] and face recognition [105].
TTF framework. Regarding TTF methods, PromptNet [106] trains a diffusion-based network that encodes the input image and noisy latent to a word embedding. To alleviate the overfitting problem, the noises predicted by the word embedding and context description are balanced through fusion sampling in classifier-free guidance. Additionally, Celeb Basis [65] provides a novel idea that the personalized ID can be viewed as the composition of celebrity’s face, which has been learned by the pre-trained diffusion model. Based on this hypothesis, a simple MLP is optimized at test time to transform face features into the weighting of different celebrity name embeddings.
PTA framework. Thanks to the abundance of available datasets featuring the same individual in various contexts, which provide valuable data for pre-training PTA methods, the number of works in PTA frameworks is rapidly increasing. Face0 [107] crops the face region to extract refined embeddings and concatenates them with text features. During the sampling phase, the output of classifier-free guidance is replaced by a weighted combination of the noise patterns predicted by face-only embedding, text-only embedding, and concatenated face-text embedding. The $\mathcal{W}+$ Adapter [53] constructs a mapping network and residual cross-attention modules to transform the facial features from the StyleGAN [108] $\mathcal{W}+$ space into the text embedding space of Stable Diffusion. FaceStudio [109] adapts the cross-attention layer to support hybrid guidance including stylized images, facial images, and textual prompts. Moreover, PhotoMaker [46] constructs a high-quality dataset through a meticulous data collection and filtering pipeline. They use a two-layer MLP to fuse ID features and class embeddings for an overall representation of human portrait. PortraitBooth [50] also employs a simple MLP, which fuses the text condition and shallow features of a pre-trained face recognition model. To ensure expression manipulation and facial fidelity, they add another expression token and incorporate the identity preservation loss and mask-based cross-attention loss. InstantID [6] additionally introduces a variant of ControlNet that takes facial landmarks as input, providing stronger guiding signals compared to the methods that solely rely on attention fusion.
@@ 5.4 Multiple Subject Composition
TTF framework. This task presents a challenge for the TTF methods, particularly in how to integrate the parameters within the same module which are separately fine-tuned for individual SoI. Some works focus on the one-for-one generation following a fusion mechanism. For instance, Custom Diffusion [20] proposes a constrained optimization method to merge the cross-attention key-value projection weights with the goal of maximizing reconstruction performance for each subject. Mix-of-Show [25] fuses the LoRA [110] weights with the same optimization objective. StyleDrop [24] dynamically summarizes noise predictions from each personalized diffusion model. In OMG [27], the latent predicted by each LoRA-tuned model is spatially composited using the subject mask. Joint training is another strategy to cover all expected subjects. SVDiff [22] employs a data augmentation method called Cut-Mix to compose several subjects together and applies a location loss to regularize attention maps, ensuring alignment between each subject and its corresponding token. Similar strategies are found in other works [20, 65] which train a single model by reconstructing the appearance of every SoI. There are advanced control mechanisms designed to manage multiple subjects. Cones [111] proposes to find a small cluster of neurons that preserve the most information about SoI. The neurons belonging to different SoI will be simultaneously activated to generate the combination. Compositional Inversion [87] introduces spatial region assignment to different subjects to improve the composition success rate.
PTA framework. For PTA frameworks, multi-subject generation is achieved through specialized architectural designs. Fastcomposer [5], Subject-Diffusion [63], and $\lambda$-eclipse [112] place each subject feature in its corresponding placeholder in the text embedding, ensuring a seamless and efficient combination. CustomNet [113] and MIGC [114] train a PTA network that supports location control for each subject. SSR-Encoder [55] implements an encoder to selectively preserve the desired subject feature and a cross-attention module to support multi-subject feature fusion.
@@ 5.5 High-level Semantic Personalization
TTF framework. For now, all researches in this field are based on the TTF framework. ReVersion [115] intends to invert object relations from references. Specifically, they use a contrastive loss to guide the optimization of the token embedding towards specific clusters of Part-of-Speech tags, such as prepositions, nouns, and verbs. Meanwhile, they also increase the likelihood of adding noise at larger timesteps during the training process to emphasize the extraction of high-level semantic features. Lego [116] focuses on more general concepts, such as adjectives, which are frequently intertwined with the subject appearance. The concept can be learned from a contrastive loss applied to the dataset comprising clean subject images and images that embody the desired adjectives. Moreover, ADI [117] aims to learn the action-specific identifier from the references. To ensure the inversion only focuses on the desired action, ADI extracts gradient invariance from a constructed triplet sample and applies a threshold to mask out the irrelevant feature channels.
@@ 6.1 Personalization on Extra Conditions
Recent personalization tasks tend to include additional conditions for diverse content customization. One common application is to customize the subject into a fixed source image. For example, PhotoSwap [118] introduces a new task that replaces the subject in the source image with the SoI from the reference image. To address this requirement, they first fine-tunes a diffusion model on the references to obtain a personalized model. To preserve the background of the source image, they initialize the noise with DDIM inversion [119] and replace the intermediate feature maps with those derived from source image generation during inference time. Later, MagiCapture [53] broadens the scope to face customization. Another similar application can be found in Virtual Try-on, which aims to fit selected clothing onto a target person. The complexities of this task have been thoroughly analyzed in another survey [120].
Additional conditions in personalization tasks may include adjusting the layout [69], transforming sketches [121], controlling viewpoint [113, 122], or modifying poses [6]. Each of these conditions presents unique challenges and requires specialized approaches to integrate these elements seamlessly into the personalized content.
@@ 6.2 Personalized Video Generation
With the rising popularity of video generation [123], video personalization has also begun to attract attention. In video personalization, the SoI can be classified into three distinct categories: appearance, motion, and the combination of both appearance and motion.
Appearance-based video personalization. This task focuses on transferring subject appearance from static images to video sequences. The standard TTF pipeline utilizes reference images as appearance anchors and fine-tunes video diffusion models (VDM) for temporal synthesis. The process involves leveraging sophisticated methods from 2D personalization, such as parameter-efficient fine-tuning [81], data augmentation [80, 81], and attention manipulation [124, 54, 81]. Additionally, several studies [125, 124, 54, 126] have explored the PTA framework. These diffusion models are specifically tailored to synthesize videos based on the image references.
Motion-based video personalization. In this task, the reference input switches to a video clip containing a consistent action. A common approach is to fine-tune the video diffusion model by reconstructing the action clip [127, 128, 129, 130, 131, 132, 133]. However, distinguishing between appearance and motion within the reference video can be challenging. To solve this problem, SAVE [128] applies appearance learning to ensure that appearance is excluded from the motion learning phase. Additionally, VMC [129] removes the background information during training prompt construction.
Appearance and motion personalization. When integrating both subject appearance and motion, innovative methods are employed to address the complexities of learning both aspects simultaneously. MotionDirector [134] utilizes spatial and temporal losses to facilitate learning across these dimensions. Another approach, DreamVideo [135], incorporates residual features from a randomly selected frame to emphasize subject information. This technique enables the fine-tuned module to primarily focus on learning motion dynamics.
@@ 6.3 Personalized 3D generation
Personalized 3D Generation refers to the process of creating customized 3D models or scenes based on 2D SoI images. Basically, the pipeline begins by fine-tuning a 2D diffusion model using TTF methods. This tuned model then utilize Score Distillation Sampling (SDS) [136] to train a 3D Neural Radiance Field (NeRF) model [137] for each specific prompt [138, 139]. Building on this foundation, several methods have been developed to improve the workflow. DreamBooth3D [140] structures the process into three phases: initializing and optimizing a NeRF from a DreamBooth model, rendering multi-view images, and fine-tuning a secondary DreamBooth for the final 3D NeRF refinement. Consist3D [141] enhances text embeddings by training two distinct tokens, a semantic and a geometric token, during 3D model optimization. TextureDreamer [142] focuses on extracting texture maps from optimized spatially-varying bidirectional reflectance distribution (BRDF) fields for rendering texture on wide-range 3D subjects.
Additionally, advancements extend to 3D avatar rendering and dynamic scenes. Animate124 [143] and Dream-in-4D [144] integrate video diffusion for 4D dynamic scene support within the 3D optimization process. In avatar rendering, PAS [145] generates 3D body poses configurable by avatar settings, StyleAvatar3D [146] facilitates 3D avatar generation based on images, and AvatarBooth [147] employs dual fine-tuned diffusion models for separate face and body generation.
@@ 6.4 Attack and Defense
This rapid advancement raises concerns about the ethical implications of PCS, particularly in areas such as misinformation, privacy violations, and the creation of deepfakes. There is an increased risk that individuals or organizations may exploit them to produce misleading content or manipulate public perception. To mitigate this, Anti-DreamBooth [148] aims to add a subtle noise perturbation to the references so that any personalized model trained on these samples only produces terrible results. The basic idea is to maximize the reconstruction loss of the surrogate model. Additionally, [149] suggests to predefine a collection of trigger words and meaningless images. These data are paired and incorporated during the training phase. Once the trigger words are encountered, the synthesized image will be intentionally altered for safeguarding.
@@ 6.5 Other Emerging Directions
Several works are exploring different personalization extensions. For example, SVG personalization is introduced by [150], in which a parameter-efficient fine-tuning method is applied to create SVGs. After first-step generation, the SVGs are refined through a process that includes semantic alignment and a dual optimization approach, which utilizes both image-level and vector-level losses to enhance the final output. Another application, 360-degree panorama customization [151], is also emerging as a potential tool for personalization in the digital imaging realm.
@@ 7.1 Evaluation Dataset
DreamBench [4] serves as the primary evaluation benchmark for DreamBooth [4], containing 30 diverse subjects (e.g., backpacks, animals, vehicles, and toys) with 25 unique prompts per subject.
DreamBench-v2 [52] expands the evaluation scope of DreamBench by adding 220 test prompts for each subject.
Custom-10 [20] used in Custom Diffusion [20] evaluates 10 subjects, each with 20 specific test prompts, and includes tests for multi-subject composition with 5 pairs of subjects and 8 prompts for each pair.
Custom-101 [20] is the latest released dataset by the authors of Custom Diffusion [20], which comprises 101 subjects to provide a broader scope of evaluation.
Stellar [35] specifically targets human-centric evaluation, featuring 20,000 prompts on 400 human identities.
@@ 7.2 Evaluation Metrics
CLIP-T measures semantic alignment using the cosine similarity between CLIP [42] embeddings of generated images and their text prompts.
ImageReward [152], HPS Score (v1/v2) [153, 154], and PickScore [155] employ learned models trained on human judgments to better correlate with perceptual quality to reflect human preference better.
@@
DINO-I [156] provides a complementary assessment using DINO’s instance-aware features, particularly effective for object-level similarity.
Fréchet Inception Distance (FID) [157] quantifies the statistical similarity between generated and real image distributions through Inception-V3 [158].
In addition to these commonly adopted metrics, there are some discussions of specialized metrics for PCS system evaluation: LyCORIS [36] introduces a 5-dimensional assessment covering fidelity, controllability, diversity, base model preservation, and image quality. Stellar [35] develops six human-centric metrics including soft-penalized CLIP text score, Identity Preservation Score, Attribute Preservation Score, Stability of Identity Score, Grounding Objects Accuracy, and Relation Fidelity Score. These metrics ensure a structured and detailed evaluation of personalized models. This evolving metric landscape reflects the growing sophistication of PCS systems, with optimal evaluation typically requiring combinations of general and task-specific measures.

### v5

@@ 5.2 Personalized Style Generation
TTF framework. In this field, StyleDrop [5sohn2024styledrop] leverages adapter tuning to efficiently capture the style from a single reference image. This method demonstrates the effectiveness through iterative training, utilizing synthesized images refined by feedback mechanisms, like human evaluations and CLIP scores. This approach not only enhances style learning but also ensures that the generated styles align closely with human aesthetic judgments. Later, GAL [151zhang2024generative] proposes an uncertainty-based evaluation strategy to filter high-quality synthetic-style data and uses a weighted schema to balance the contribution of the additional samples and the original reference. Furthermore, StyleAligned [98hertz2023style] focuses on maintaining stylistic consistency across a batch of images. This is achieved by using the first image as a reference, which acts as an additional key and value in the self-attention layers, ensuring that all subsequent images in the batch adhere to the same stylistic guidelines. Style-friendly [211choi2024style] introduces a novel diffusion model fine-tuning approach that enhances personalized artistic style generation by adaptively biasing noise sampling toward higher noise levels, where stylistic features emerge.
PTA framework. For the PTA framework, StyleAdapter [26wang2023styleadapter] employs a dual-path cross-attention mechanism within the PTA framework. This model introduces a specialized embedding module designed to extract and integrate global features from multiple style references. Diptych Prompting [209shin2024large] utilizes an inpainting mechanism to draw another image with the same style of the reference part.
@@ 5.3 Personalized Face Generation
Personalized face generation aims to generate diverse identity images that adhere to text prompt specifications, utilizing only a few initial face images. Compared to general object personalization, the scope is narrowed to a specific class, humans. An obvious benefit is that one can readily leverage large-scale human-centric datasets [zheng2022general, yang2020facescape, cao2018vggface2] and utilize pre-trained models in well-developed areas, like face landmark detection [wu2019facial] and face recognition [adjabi2020past].
TTF framework. Regarding TTF methods, PromptNet [55zhou2023enhancing] trains a diffusion-based network that encodes the input image and noisy latent to a word embedding. To alleviate the overfitting problem, the noises predicted by the word embedding and context description are balanced through fusion sampling in classifier-free guidance. Additionally, Celeb Basis [60yuan2023inserting] provides a novel idea that the personalized ID can be viewed as the composition of celebrity’s face, which has been learned by the pre-trained diffusion model. Based on this hypothesis, a simple MLP is optimized at test time to transform face features into the weighting of different celebrity name embeddings.
PTA framework. Thanks to the abundance of available datasets featuring the same individual in various contexts, which provide valuable data for pre-training PTA methods, the number of works in PTA frameworks is rapidly increasing. Face0 [88valevski2023face0] crops the face region to extract refined embeddings and concatenates them with text features. During the sampling phase, the output of classifier-free guidance is replaced by a weighted combination of the noise patterns predicted by face-only embedding, text-only embedding, and concatenated face-text embedding. The $\mathcal{W}+$ Adapter [91li2023stylegan] constructs a mapping network and residual cross-attention modules to transform the facial features from the StyleGAN [karras2019style] $\mathcal{W}+$ space into the text embedding space of Stable Diffusion. FaceStudio [31yan2023facestudio] adapts the cross-attention layer to support hybrid guidance including stylized images, facial images, and textual prompts. Moreover, PhotoMaker [28li2023photomaker] constructs a high-quality dataset through a meticulous data collection and filtering pipeline. They use a two-layer MLP to fuse ID features and class embeddings for an overall representation of human portrait. PortraitBooth [96peng2023portraitbooth] also employs a simple MLP, which fuses the text condition and shallow features of a pre-trained face recognition model. To ensure expression manipulation and facial fidelity, they add another expression token and incorporate the identity preservation loss and mask-based cross-attention loss. InstantID [37wang2024instantid] additionally introduces a variant of ControlNet that takes facial landmarks as input, providing stronger guiding signals compared to the methods that solely rely on attention fusion.
@@ 5.4 Multiple Subject Composition
TTF framework. This task presents a challenge for the TTF methods, particularly in how to integrate the parameters within the same module which are separately fine-tuned for individual SoI. Some works focus on the one-for-one generation following a fusion mechanism. For instance, Custom Diffusion [6kumari2023multi] proposes a constrained optimization method to merge the cross-attention key-value projection weights with the goal of maximizing reconstruction performance for each subject. Mix-of-Show [46gu2024mix] fuses the LoRA [hu2021lora] weights with the same optimization objective. StyleDrop [5sohn2024styledrop] dynamically summarizes noise predictions from each personalized diffusion model. In OMG [115kong2024omg], the latent predicted by each LoRA-tuned model is spatially composited using the subject mask. Joint training is another strategy to cover all expected subjects. SVDiff [8han2023svdiff] employs a data augmentation method called Cut-Mix to compose several subjects together and applies a location loss to regularize attention maps, ensuring alignment between each subject and its corresponding token. Similar strategies are found in other works [6kumari2023multi, 60yuan2023inserting] which train a single model by reconstructing the appearance of every SoI. There are advanced control mechanisms designed to manage multiple subjects. Cones [4liu2023cones] proposes to find a small cluster of neurons that preserve the most information about SoI. The neurons belonging to different SoI will be simultaneously activated to generate the combination. Compositional Inversion [9zhang2023compositional] introduces spatial region assignment to different subjects to improve the composition success rate.
PTA framework. For PTA frameworks, multi-subject generation is achieved through specialized architectural designs. Fastcomposer [54xiao2023fastcomposer], Subject-Diffusion [16ma2023subject], and $\lambda$-eclipse [222patel2024lambda] place each subject feature in its corresponding placeholder in the text embedding, ensuring a seamless and efficient combination. CustomNet [29yuan2023customnet] and MIGC [174zhou2024migc] train a PTA network that supports location control for each subject. SSR-Encoder [221zhang2024ssr] implements an encoder to selectively preserve the desired subject feature and a cross-attention module to support multi-subject feature fusion.
@@ 5.5 High-level Semantic Personalization
TTF framework. For now, all researches in this field are based on the TTF framework. ReVersion [63huang2023reversion] intends to invert object relations from references. Specifically, they use a contrastive loss to guide the optimization of the token embedding towards specific clusters of Part-of-Speech tags, such as prepositions, nouns, and verbs. Meanwhile, they also increase the likelihood of adding noise at larger timesteps during the training process to emphasize the extraction of high-level semantic features. Lego [45motamed2023lego] focuses on more general concepts, such as adjectives, which are frequently intertwined with the subject appearance. The concept can be learned from a contrastive loss applied to the dataset comprising clean subject images and images that embody the desired adjectives. Moreover, ADI [43huang2023learning] aims to learn the action-specific identifier from the references. To ensure the inversion only focuses on the desired action, ADI extracts gradient invariance from a constructed triplet sample and applies a threshold to mask out the irrelevant feature channels.
@@ 6.1 Personalization on Extra Conditions
Recent personalization tasks tend to include additional conditions for diverse content customization. One common application is to customize the subject into a fixed source image. For example, PhotoSwap [15gu2024photoswap] introduces a new task that replaces the subject in the source image with the SoI from the reference image. To address this requirement, they first fine-tunes a diffusion model on the references to obtain a personalized model. To preserve the background of the source image, they initialize the noise with DDIM inversion [song2020denoising] and replace the intermediate feature maps with those derived from source image generation during inference time. Later, MagiCapture [91li2023stylegan] broadens the scope to face customization. Another similar application can be found in Virtual Try-on, which aims to fit selected clothing onto a target person. The complexities of this task have been thoroughly analyzed in another survey [islam2024deep].
Additional conditions in personalization tasks may include adjusting the layout [152chen2024training], transforming sketches [94chen2023democaricature], controlling viewpoint [29yuan2023customnet, 163kumari2024customizing], or modifying poses [37wang2024instantid]. Each of these conditions presents unique challenges and requires specialized approaches to integrate these elements seamlessly into the personalized content.
@@ 6.2 Personalized Video Generation
With the rising popularity of video generation [xing2024survey], video personalization has also begun to attract attention. In video personalization, the SoI can be classified into three distinct categories: appearance, motion, and the combination of both appearance and motion.
Appearance-based video personalization. This task focuses on transferring subject appearance from static images to video sequences. The standard TTF pipeline utilizes reference images as appearance anchors and fine-tunes video diffusion models (VDM) for temporal synthesis. The process involves leveraging sophisticated methods from 2D personalization, such as parameter-efficient fine-tuning [145wang2024customvideo], data augmentation [138chen2023videodreamer, 145wang2024customvideo], and attention manipulation [83jiang2023videobooth, 212liu2024stylecrafter, 145wang2024customvideo]. Additionally, several studies [82zhao2023videoassembler, 83jiang2023videobooth, 212liu2024stylecrafter, 105feng2023dreamoving] have explored the PTA framework. These diffusion models are specifically tailored to synthesize videos based on the image references.
Motion-based video personalization. In this task, the reference input switches to a video clip containing a consistent action. A common approach is to fine-tune the video diffusion model by reconstructing the action clip [150wu2023tune, 103song2023save, 140jeong2023vmc, 148esser2023structure, 136he2023animate, 102wu2023lamp, 104materzynska2023customizing]. However, distinguishing between appearance and motion within the reference video can be challenging. To solve this problem, SAVE [103song2023save] applies appearance learning to ensure that appearance is excluded from the motion learning phase. Additionally, VMC [140jeong2023vmc] removes the background information during training prompt construction.
Appearance and motion personalization. When integrating both subject appearance and motion, innovative methods are employed to address the complexities of learning both aspects simultaneously. MotionDirector [101zhao2023motiondirector] utilizes spatial and temporal losses to facilitate learning across these dimensions. Another approach, DreamVideo [142wei2023dreamvideo], incorporates residual features from a randomly selected frame to emphasize subject information. This technique enables the fine-tuned module to primarily focus on learning motion dynamics.
@@ 6.3 Personalized 3D generation
Personalized 3D Generation refers to the process of creating customized 3D models or scenes based on 2D SoI images. Basically, the pipeline begins by fine-tuning a 2D diffusion model using TTF methods. This tuned model then utilize Score Distillation Sampling (SDS) [poole2022dreamfusion] to train a 3D Neural Radiance Field (NeRF) model [mildenhall2021nerf] for each specific prompt [120lin2023magic3d, 126shi2023mvdream]. Building on this foundation, several methods have been developed to improve the workflow. DreamBooth3D [121raj2023dreambooth3d] structures the process into three phases: initializing and optimizing a NeRF from a DreamBooth model, rendering multi-view images, and fine-tuning a secondary DreamBooth for the final 3D NeRF refinement. Consist3D [127ouyang2023chasing] enhances text embeddings by training two distinct tokens, a semantic and a geometric token, during 3D model optimization. TextureDreamer [133yeh2024texturedreamer] focuses on extracting texture maps from optimized spatially-varying bidirectional reflectance distribution (BRDF) fields for rendering texture on wide-range 3D subjects.
Additionally, advancements extend to 3D avatar rendering and dynamic scenes. Animate124 [128zhao2023animate124] and Dream-in-4D [130zheng2023unified] integrate video diffusion for 4D dynamic scene support within the 3D optimization process. In avatar rendering, PAS [122azadi2023text] generates 3D body poses configurable by avatar settings, StyleAvatar3D [123zhang2023styleavatar3d] facilitates 3D avatar generation based on images, and AvatarBooth [124zeng2023avatarbooth] employs dual fine-tuned diffusion models for separate face and body generation.
@@ 6.4 Attack and Defense
This rapid advancement raises concerns about the ethical implications of PCS, particularly in areas such as misinformation, privacy violations, and the creation of deepfakes. There is an increased risk that individuals or organizations may exploit them to produce misleading content or manipulate public perception. To mitigate this, Anti-DreamBooth [13van2023anti] aims to add a subtle noise perturbation to the references so that any personalized model trained on these samples only produces terrible results. The basic idea is to maximize the reconstruction loss of the surrogate model. Additionally, [50zhang2023backdooring] suggests to predefine a collection of trigger words and meaningless images. These data are paired and incorporated during the training phase. Once the trigger words are encountered, the synthesized image will be intentionally altered for safeguarding.
@@ 6.5 Other Emerging Directions
Several works are exploring different personalization extensions. For example, SVG personalization is introduced by [36zhang2023text], in which a parameter-efficient fine-tuning method is applied to create SVGs. After first-step generation, the SVGs are refined through a process that includes semantic alignment and a dual optimization approach, which utilizes both image-level and vector-level losses to enhance the final output. Another application, 360-degree panorama customization [25wang2024customizing], is also emerging as a potential tool for personalization in the digital imaging realm.
@@ 7.1 Evaluation Dataset
DreamBench [2ruiz2023dreambooth] serves as the primary evaluation benchmark for DreamBooth [2ruiz2023dreambooth], containing 30 diverse subjects (e.g., backpacks, animals, vehicles, and toys) with 25 unique prompts per subject.
DreamBench-v2 [11chen2024subject] expands the evaluation scope of DreamBench by adding 220 test prompts for each subject.
Custom-10 [6kumari2023multi] used in Custom Diffusion [6kumari2023multi] evaluates 10 subjects, each with 20 specific test prompts, and includes tests for multi-subject composition with 5 pairs of subjects and 8 prompts for each pair.
Custom-101 [6kumari2023multi] is the latest released dataset by the authors of Custom Diffusion [6kumari2023multi], which comprises 101 subjects to provide a broader scope of evaluation.
Stellar [95achlioptas2023stellar] specifically targets human-centric evaluation, featuring 20,000 prompts on 400 human identities.
@@ 7.2 Evaluation Metrics
CLIP-T measures semantic alignment using the cosine similarity between CLIP [radford2021learning] embeddings of generated images and their text prompts.
ImageReward [xu2023imagereward], HPS Score (v1/v2) [wu2023human, wu2023humanv2], and PickScore [kirstain2023pick] employ learned models trained on human judgments to better correlate with perceptual quality to reflect human preference better.
@@
DINO-I [caron2021emerging] provides a complementary assessment using DINO’s instance-aware features, particularly effective for object-level similarity.
Fréchet Inception Distance (FID) [FIDheusel2017gans] quantifies the statistical similarity between generated and real image distributions through Inception-V3 [inception].
In addition to these commonly adopted metrics, there are some discussions of specialized metrics for PCS system evaluation: LyCORIS [76yeh2023navigating] introduces a 5-dimensional assessment covering fidelity, controllability, diversity, base model preservation, and image quality. Stellar [95achlioptas2023stellar] develops six human-centric metrics including soft-penalized CLIP text score, Identity Preservation Score, Attribute Preservation Score, Stability of Identity Score, Grounding Objects Accuracy, and Relation Fidelity Score. These metrics ensure a structured and detailed evaluation of personalized models. This evolving metric landscape reflects the growing sophistication of PCS systems, with optimal evaluation typically requiring combinations of general and task-specific measures.